OpenStudy (anonymous):

Evaluate log wit base of 2 1 over 64.

OpenStudy (solomonzelman):

\[\log_2\frac{1}{64}\] this ?

OpenStudy (anonymous):

yes. would it be -6?

OpenStudy (solomonzelman):

\[\log_2(\frac{1}{2})^{~6}\]\[6~\log_2(\frac{1}{2})\]\[6~(\log_21-\log_22)\]\[6(\log_21-1)\]\[6\log_21-6\]